OpenStudy (anonymous):

Given f(x) =5x+1/2solve for f^1(8) 3 7 10 20

OpenStudy (austinl):

Is that \(f^{-1}(8)\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

the equation is\[f ^{-1}(8)=5x+1\div2\]

OpenStudy (anonymous):

Ugh! Austin you always beat me to them XD

OpenStudy (austinl):

\(\displaystyle f^{-1}(x)=5x+\frac{1}{2}\) Or \(\displaystyle f^{-1}(x)=\frac{5x+1}{2}\)

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

wait it not negetive power it is positive one\[f^{1}\]

OpenStudy (austinl):

Actually, that would be \(\displaystyle f(x)=\frac{5x+1}{2}\) We need to find \(f^{-1}(x)\) How we would do this is as follows. Set f(x)=y so now we have, \(\displaystyle y=\frac{5x+1}{2}\) Now we need to solve for x. \(\displaystyle y=\frac{5x+1}{2}\) \(2y=5x+1\) \(2y-1=5x\) \(\displaystyle x=\frac{2y-1}{5}\) Now we switch the x and y. \(\displaystyle y=\frac{2x-1}{5}\) \(\displaystyle f^{-1}(x)=\frac{2x-1}{5}\) Now we need to find \(f^{-1}(8)\) Which all we would do is simply replace all instances of x with 8 in the function. Do you think you can do that?

OpenStudy (anonymous):

yea but change the f^-1 to f^1 it glitch when i typed it

OpenStudy (austinl):

That makes absolutely no sense... \(f^{1}(x)=f(x)\)....

OpenStudy (anonymous):

idk either but tht is how it is on my test

OpenStudy (austinl):

wait.... this is a test?

OpenStudy (anonymous):

it a redo.i did the question but i got it wrong the first time.it confused me so she said i could ask for help on tht question

OpenStudy (austinl):

Ok then... well... just plug in 8 for all instances of x in f(x)

OpenStudy (anonymous):

ok i did and i got [40+1 \div2\]

OpenStudy (austinl):

\(\displaystyle f(8)=\frac{5(8)+1}{2}=\frac{41}{2}\) That should be your answer with the weird notation and all.

OpenStudy (anonymous):

oh i forgot to add the one.ok then would i divide?

OpenStudy (austinl):

You can leave it like that, or you can make it an improper fraction. But that should be acceptable as is.

OpenStudy (anonymous):

ok thank you

OpenStudy (austinl):

No problem!